72% of adults in a certain country believe that life on other planets is plausible. You randomly select five adults and ask them whether they believe that life on other planets is plausible. The random variable represents the number of adults who believe that life on other planets is plausible.

Find the mean, variance, and standard deviation of the binomial distribution for the random variable.

Interpret the results.

Find the mean of the binomial distribution.

mu =    (Round to two decimal places as needed.)
Answer #1

Here the probability of success i.e. p = 72/100= 0.72 and q = 1-0.72 = 0.28

Given : n=5

For Binomial distribution:

Mean : np = 5*0.72 = 3.6

Variance : npq = 5*0.72*0.28 = 1.008

Standard deviation : square root of npq = 1.004